OpenStudy (doodleduck8):

Solve the equation for y: 4x + 3y = -12

OpenStudy (abdumu16):

y=1/3*(-12-4x)

OpenStudy (3mar):

If you need to solve for y, the idea is to separate y in one side and all other terms in the other side! That is accomplished as follows: \[\Large \bf\color{#F29ECF} {4x + 3y = -12}\] send \(4x\) to the other side to be with \(-12\), or subtract \(4x\) from both sides: \[\Large \bf\color{#CCADD6} {4x -4x+ 3y = -12-4x}\] SIMPLIFY:: \[\Large \bf\color{#B2B8DB} {3y = -4x-12}\] Divide both sides be 3 to get the coefficient of \(y\) unity: \[\Large \bf\color{#8CC7E3} {\frac{ 3y }{ 3 } = \frac{ -4x-12 }{ 3 }}\] SIMPLIFY:: \[\Large \bf\color{#73D1E8} {y= -\frac{ 4}{ 3 }x-\frac{ 12 }{ 3 }}\] \[\Huge \bf\color{#4DE0F0} {y= -\frac{ 4}{ 3 }x-4}\] That is the from of the slope-intercept of a line (solved for y) I hope I helped you.

OpenStudy (studybuddy26):

You first need to get the y alone .

OpenStudy (studybuddy26):

4x + 3y = -12 3y = -12 - 4x Y= (-12-4x )/3

OpenStudy (studybuddy26):

Now just finish it off . :)

OpenStudy (doodleduck8):

Thanks, that really helped

OpenStudy (doodleduck8):

@studybuddy26

OpenStudy (studybuddy26):

Your welcome

OpenStudy (3mar):

Thank you for the medal, @princeharryyy and @narissa! I appreciate!

OpenStudy (princeharryyy):

:)